OpenStudy (anonymous):

Three consecutive integers with the sum of 42

OpenStudy (anonymous):

@oldrin.bataku

OpenStudy (anonymous):

What is 42/3? It's 14 Now you have to have numbers that are equal to 14(3) 13+14+15 This is very similar to the last question I showed you. Is the concept sinking in?

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

so 14 is one of them..

OpenStudy (texaschic101):

3 consecutive integers....that means one right after the other Let x be the first integer, and let x + 1 be the second integer, and let x + 2 be the third integer. x + x+ 1 + x + 2 = 42 -- combine like terms 3x + 3 = 42 -- subtract 3 from both sides 3x = 42 - 3 3x = 39 -- divide by 3 x = 13

OpenStudy (anonymous):

we're going to use the same method as before. let \(n\) be our middle integer, so the first and third integers are \(n-1\) and \(n+1\) as a I mentioned before. when we say the sum is \(42\) we mean that if we add \(n-1,n,n+1\) it should result in \(42\) i.e.$$(n-1)+n+(n+1)=42\\\dots$$

OpenStudy (anonymous):

solve for \(n\) to determine the middle number. the first number will then be the one right before it (\(n-1\)) and the third will be the one right after it (\(n+1\))

OpenStudy (anonymous):

Oh I see. oldrin has the right idea Look at the formula two messages up. 14 is your "n" Plug 14 into the formula and see if it works. Remember you got 14 because 42/3 = 14

OpenStudy (anonymous):

so its 13 14 n 15

OpenStudy (anonymous):

but how would it be if it were to even integers

OpenStudy (anonymous):

Correct :)

OpenStudy (anonymous):

14 and 12 or 14 and 16

OpenStudy (texaschic101):

it does not say even integers, it says consecutive integers

OpenStudy (anonymous):

Absolutely. As long as the the middle number is 14 The middle number of 13 and 15 is 15 The middle number of 12 and 16 is 14

OpenStudy (anonymous):

Sorry I mean The middle number of 13 and 15 is 14 And yes texaschic is correct. The question asks consecutively.

OpenStudy (anonymous):

if they asked for only even integers, then instead of using \(n-1,n,n+1\) we would use \(n-2,n,n+2\) -- the process is very similar. the only difference is that consecutive integers are only \(1\) apart whereas even integers are \(2\) apart

OpenStudy (texaschic101):

or n , n + 2, and n + 4

OpenStudy (anonymous):

what about for 159

OpenStudy (anonymous):

52 53 and 54?

OpenStudy (anonymous):

@texaschic101 sure, if you pick \(n\) to represent the first number -- but it's more natural to consider the middle number due to the symmetry of the problem. observe that when you sum \(n,n+2,n+4\) you get \(3n+6\) whereas summing \(n-2,n,n+2\) gives just \(3n\)

OpenStudy (anonymous):

nvm it would be 79 and 80

OpenStudy (texaschic101):

n + n + 1 + n + 2 = 159 3n + 3 = 159 3n = 159 - 3 3n = 156 n = 52 n + 1 = 53 n + 2 = 54